1. The nurse is administering amoxicillin to her patient with a sinus infection. Which is the best way for her to insure that she is giving it to the right patient?

A) A Call the patient by name

B)Check the patient’s wristband

C)Read the name of the patient on the patient’s door

D) Check the patient’s room number on the unit census list

2. Mrs. Razo has been diagnosed with herpes zoster and is prescribed acyclovir (Zovirax) 800 mg, 5 times per day × 7 days. The pharmacy has acyclovir (Zovirax) available as 200-mg capsules. What is the total number of capsules Mrs. Rahnned should receive in her 7-day prescription?

3. Doctor’s Order an antibiotic for Mr. Smith's wound infection. Doctor's order: Rocephin 0.5 grams in 250 mL of D5W to infuse IV over 45 minutes; Drop Factor: 12gtt/min. How many gtt/ min will you regulate the IVPB? (Remember to round to the nearest whole number)

A) 30gtts/min

B) 67gtts/min

C) 87gtts/min

D) 6gtts/min

4. Mr. Smith takes in 1 and LaTeX: \frac{1}{2} 1 2 cup of water, 6 oz. of coffee and 2 tablespoons of broth (which you will add in) How many mL will you document?

5. When you are monitoring your client who is now started on an intravenous antibiotic for an infection, you notice that the client is exhibiting signs of anaphylaxis. What is your first priority intervention?

A) Slow down the intravenous flow

B) Stop the intravenous flow

C) Notify the doctor

D) Begin CPR
Answer #1

Q.1. The best way to idetify the correct patient before giving the medication is to check the patient's wrist band, which provides the proper medication administration record of that patient as per the set therapeutic plan.Hence option B is correct.

2. Prescribed acyclovir (Zovirax) = 800 mg, 5 times per day × 7 days.

Available acyclovir (Zovirax) = 200-mg capsules.

Required capsules per dose =800/200= 4 capsules

Per day 5 times means 5x4=20 capsules per day

For 7 days = 20 x7=140 capsules.

3. IV infusion rate (gtts/min) = 250 ml/(45 min) X 12 gtts/mL

                                           = 66.67= 67 gtts/min

Hence option B is correct.

4. Smith takes1 and half cup of water, 6 oz. of coffee,  2 tablespoons of broth

Total ml =?

1 US cup of water =236.59 ml

Then one and half cup of water= 236.59 + 118.295= 354.88 mL

6 oz. of water= 177 ml

2 table spoons= 29.57

Total mL= 354.88 + 177 + 29.57= 561.45 mL.

5. Option B is correct as there is need to slow down the intravenous flow as it may be due to anaphylaxis reaction of the antibiotics which is administered.